OpenStudy (aaronandyson):

A can do a piece of work in x day and B cando the same work in x+6 days.If both working together can do it in 15 days;calculate x @sepeario

OpenStudy (aaronandyson):

@Sepeario

OpenStudy (sepeario):

Sorry I stuffed up. \[\frac{ A }{ x }+\frac{ B }{ x+6 }=15\]

OpenStudy (sepeario):

Can you further expand/evaluate that?

OpenStudy (anonymous):

a=x b=x+6 a+b=16 sub a and b into a+b=16 x+x+6=16 2x=10 x=5

OpenStudy (aaronandyson):

I'm confused

OpenStudy (anonymous):

understand now?

OpenStudy (aaronandyson):

It's 15 days not 16 :S

OpenStudy (anonymous):

oh whoops

OpenStudy (anonymous):

a=x b=x+6 a+b=15 sub a and b into a+b=15 x+x+6=15 2x=9 x=9/2 or 4.5

OpenStudy (anonymous):

now you get it :)

OpenStudy (anonymous):

so A can do a piece of work in 4.5 days B can do a piece of work in 10.5 days and together they can do a piece of work in 4.5days

OpenStudy (sepeario):

Very well explained :)

OpenStudy (anonymous):

thanks @Sepeario

OpenStudy (aaronandyson):

This method is simple but the ans is 24 :/

OpenStudy (anonymous):

thats weird?

OpenStudy (anonymous):

can you check if your question you posted was right

OpenStudy (aaronandyson):

I have posted the question word for word .

OpenStudy (aaronandyson):

@Sepeario

OpenStudy (aaronandyson):

Nevertheless ,I'll ask my teacher xD Can u help me with 3 more question?

OpenStudy (anonymous):

okay